OpenStudy (anonymous):

If a stone is tossed from the top of a 330 meter building, the height of the stone as a function of time is given by h(t) = -9.8t2 – 10t + 330, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer. Please show your work.

OpenStudy (anonymous):

just solve \[-9,8t^2-10t+330=0,t \ge 0\] Use the quadratic formula.

OpenStudy (anonymous):

I tried

OpenStudy (anonymous):

I cant get any method to work

OpenStudy (anonymous):

The quadratic formula is \[\left[-(b) \pm \sqrt{((b) ^{2}-4(a)(c))}\right]/(2(a))\]so your equation looks like this -9.8t^2-10t+330 and it is in the form at^2-bt+c, so a is -9.8, b is -10 and c is 330. Plug those into the equation. \[\left[-(-10) \pm \sqrt{((-10) ^{2}-4(-9.8)(330))}\right]/(2(-9.8))\] if you simplify that you will get: \[(10\pm \ \sqrt{13036} )/-18.6\] That can be simplified to \[(10\pm2\sqrt{3259})/-18.6\] we can eliminate the case where it will be negative (if we use the positive sign it will maek the qhole thing negative and we can't have a negative time). So the negative case is \[(10-2\sqrt{3259})/-18.6 \] which is approximately 5.6 seconds.

OpenStudy (anonymous):

Oops. The denominator should be -19.6 not -18.6 so the positive answer is x = 5.32 secs to the nearest hundredth.

OpenStudy (anonymous):

You're correct; my mistake.

OpenStudy (anonymous):

Is there a way to edit answers lol...